OpenStudy (anonymous):

Let R be the region in the first quadrant bounded by the y-axis and the graphs of y=4x-x^(3)+1 and y=3/4x. a)Find the area of R. b)Find the volume of the solid generated when R is revovled about the x-axis. c)Write an expression involving one or more integrals that gives the perimeter of R.Do not evaluate

OpenStudy (amistre64):

function - function determines height

OpenStudy (amistre64):

\[\int f(x)-\int g(x)\] from 0 to their intersection i spose; unless i missed an ending interval

OpenStudy (anonymous):

They intersect where 4x - x^3 +1 = 3x/4 so you could solve a cubic - or guess! So x^3 - 13x/4 -1 = 0 Hmmmm